# A novel measurement of the strong-phase difference between $D^{0}\to K^{-}{{\pi}^{+}}$ and $\bar{D}^{0}\to K^{-}{{\pi}^{+}}$ decays using $C$-even and $C$-odd quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs

### v1

@@ III Selection Requirements
Finally, two techniques can be applied to reconstruct ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced in $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays. The first proceeds by reconstructing a single $D^{*0}\to{{\pi}^{0}}D^{0}$ candidate. However, due to the low momenta of the photons produced in the ${{\pi}^{0}}\to\gamma\gamma$ decay, the reconstruction efficiency is too low. The second attempt examines the candidates that do not pass the requirements placed to reconstruct the other production mechanisms originating from the $D^{*}\bar{D}^{*}$ process. However, the irreducible background from the $D^{*+}D^{*-}$ process is deemed to be too large. Therefore, in this analysis, the $e^{+}e^{-}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism is not explicitly reconstructed. The small amount that remain in the sample are incorporated as signal alongside the $e^{+}e^{-}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair has the same $C$-eigenvalue. Together they are referred to as $D^{*}\bar{D}^{*}\to{\gamma\gamma/\pi^{0}\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$.
Table 2 summarizes the selection requirements applied to isolate the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs. As mentioned above, and shown in Fig. 1, the criteria imposed do not isolate pure samples, and there is non-negligible cross-feed between the production mechanisms. Table 3 displays the efficiencies determined in simulation of reconstructing a particular production mechanism given the selection requirements applied to isolate another. It should be noted that the efficiency of correctly reconstructing ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from the $e^{+}e^{-}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism with the designed selection requirements is relatively high ($\sim 25\%$), but it is reduced in Table 3 because the definition of the $D^{*}\bar{D}^{*}\to{\gamma\gamma/\pi^{0}\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanism includes the $e^{+}e^{-}\to{{\pi}^{0}}{{\pi}^{0}}D^{*}\bar{D}^{*}$ process.
@@ V.1 Overview
A measurement of the strong-phase difference between $D^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ and $\bar{D}^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decays, $\delta_{K\pi}^{D}$, is performed. This determination assumes the expected quantum-correlated behavior of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from each production mechanism that is validated in Sec. IV. The strong-phase difference $\delta_{K\pi}^{D}$ is measured using ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ Y$ decays, where $Y$ is one of the four $C\!P$ eigenstates presented in Table 4 or ${K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$, and is referred to as the ‘tag’ throughout.
@@
Inputs for the hadronic parameters of the ${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decay are necessary to extract the observables. They are the fraction of $D^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays that are in region $i$, denoted $K_{i}$, and the amplitude weighted averaged cosine (sine) difference of the strong-phase difference between $D^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ and $\bar{D}^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays, denoted $c_{i}$ ($s_{i}$). They are defined as
@@
${K_{-i}}$ $=\int_{i}|A(\bar{D}^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|^{2}\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (23)
${c_{i}}$ $=\frac{1}{\sqrt{{K_{i}}{K_{-i}}}}\int_{i}|A(D^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})||A(\bar{D}^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|\cos(\Delta\delta_{D})\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (24)
${s_{i}}$ $=\frac{1}{\sqrt{{K_{i}}{K_{-i}}}}\int_{i}|A(D^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})||A(\bar{D}^{0}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|\sin(\Delta\delta_{D})\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (25)
where $\Delta\delta_{D}$ is the strong-phase difference between $D^{0}$ and $\bar{D}^{0}$ decays to the ${K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ final state. The values used in this measurement for each are taken from Ref. [34]. With these definitions, the parameters in Eq. (4) become $r_{1}\to{r_{K\pi}^{D}}$, $\delta_{1}\to\delta_{K\pi}^{D}$, $R_{1}\to R_{K\pi}=1$, $A_{2}^{2}\to{K_{i}}$, $r_{2}^{2}A_{2}^{2}\to{K_{-i}}$, $R_{2}\cos\delta_{2}\to{c_{i}}$ and $R_{2}\sin\delta_{2}\to{s_{i}}$, such that the fraction of ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays in a region, $i$, is given by,
@@ V.2 Measuring ${{r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}}$ using $D$ decays to $C\!P$-eigenstates
Figures 8 and 9 display the fits for each $C\!P$ tag of the candidates that pass the selection criteria to isolate the $D^{*}\bar{D}\to{\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $D^{*}\bar{D}^{*}\to{\gamma\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanisms, which have the largest yields. The fitted signal yields are corrected for residual backgrounds from ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays to the correct final state, where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair originates from the $e^{+}e^{-}\to D^{*+}D^{*-}$ process, by subtracting the yield in simulation. The corrections are only applied to ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs originating from the $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ processes and in all cases are less than 7% of the fitted yields. The signal yields after the corrections are displayed in Table 7.
@@
The parameters ${r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}$ and ${r_{K\pi}^{D}}\sin\delta_{K\pi}^{D}$ are determined by simultaneously fitting the reconstructed invariant-mass distribution of both $D$ mesons in 80 categories given by the region of the ${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ phase space and the hypothesized production mechanism. In addition to the five core components discussed previously, two additional background components are included in the fit. The first is from $D^{0}{{D}^{-}}\to[{{\textit{K}}^{-}}{{\pi}^{+}}]_{D^{0}},[{K^{0}_{\mathrm{S}}}{{\pi}^{-}}{{\pi}^{0}}]_{{D}^{-}}$ decays, where the $D^{0}{{D}^{-}}$ pair originates from $e^{+}e^{-}\to{{D}^{*+}}{{D}^{-}}\to D^{0}{{\pi}^{+}}{{D}^{-}}$ production. These decays are only present in the categories which isolate ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs originating from the $e^{+}e^{-}\to D^{*}\bar{D}$ process. The selection requirements effectively remove this background in the other categories. The shape and yield of this component are fixed to those determined in simulation. The second background is from ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays, where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair originates from $e^{+}e^{-}\to D^{*+}D^{*-}$ production. The background is only present in the categories which isolate ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced by the $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ process. In the fit, this background is modeled with the same shape as signal and the yield is fixed to that determined in simulation.
@@ VI Systematic uncertainties
The cross-feed between the production mechanisms is primarily caused by missing energy due to ISR. In events with ISR the effective center-of-mass energy of the $e^{+}e^{-}$ pair is reduced. To model such events the simulation requires knowledge of the cross sections of the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $e^{+}e^{-}\to D^{*}\bar{D}$ and $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ processes at the reduced center-of-mass energies, where the relative uncertainties are at most around 20% [36]. The associated systematic uncertainty is determined by repeating the fits which determine the observables using alternate efficiency matrices which are calculated in simulation where the events with ISR are reweighted by ±$\pm 20\%$. Despite this conservative estimate the shifts in the observables are relatively small, for ${r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}$ measured with the $C\!P$ tags (${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$) it is around 10% (4%) of the statistical uncertainty, whilst for ${r_{K\pi}^{D}}\sin\delta_{K\pi}^{D}$ it is found to be negligible. This systematic uncertainty is fully correlated across all observables.

### v2

@@ III Selection Requirements
Finally, two techniques can be applied to reconstruct ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced in ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays. The first proceeds by reconstructing a single $D^{*0}\to{{\pi}^{0}}{{D}^{0}}$ candidate. However, due to the low momenta of the photons produced in the ${{\pi}^{0}}\to\gamma\gamma$ decay, the reconstruction efficiency is too low. The second attempt examines the candidates that do not pass the requirements placed to reconstruct the other production mechanisms originating from the ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ process. However, the irreducible background from the $D^{*+}D^{*-}$ process is deemed to be too large. Therefore, in this analysis, the $e^{+}e^{-}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism is not explicitly reconstructed. The small amount that remain in the sample are incorporated as signal alongside the $e^{+}e^{-}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair has the same $C$-eigenvalue. Together they are referred to as $D^{*}\bar{D}^{*}\to{\gamma\gamma/\pi^{0}\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$.
Table 2 summarizes the selection requirements applied to isolate the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs. As mentioned above, and shown in Fig. 1, the criteria imposed do not isolate pure samples, and there is non-negligible cross-feed between the production mechanisms. Table 3 displays the efficiencies determined in simulation of reconstructing a particular production mechanism given the selection requirements applied to isolate another. It should be noted that the efficiency of correctly reconstructing ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from the $e^{+}e^{-}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism with the designed selection requirements is relatively high ($\sim 25\%$), but it is reduced in Table 3 because the definition of the $D^{*}\bar{D}^{*}\to{\gamma\gamma/\pi^{0}\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanism includes the $e^{+}e^{-}\to{{\pi}^{0}}{{\pi}^{0}}{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ process.
@@ V.1 Overview
A measurement of the strong-phase difference between ${{D}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ and ${{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decays, $\delta_{K\pi}^{D}$, is performed. This determination assumes the expected quantum-correlated behavior of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from each production mechanism that is validated in Sec. IV. The strong-phase difference $\delta_{K\pi}^{D}$ is measured using ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ Y$ decays, where $Y$ is one of the four $C\!P$ eigenstates presented in Table 4 or ${K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$, and is referred to as the ‘tag’ throughout.
@@
Inputs for the hadronic parameters of the ${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decay are necessary to extract the observables. They are the fraction of ${{D}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays that are in region $i$, denoted $K_{i}$, and the amplitude weighted averaged cosine (sine) difference of the strong-phase difference between ${{D}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ and ${{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays, denoted $c_{i}$ ($s_{i}$). They are defined as
@@
${K_{-i}}$ $=\int_{i}|A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|^{2}\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (23)
${c_{i}}$ $=\frac{1}{\sqrt{{K_{i}}{K_{-i}}}}\int_{i}|A({{D}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})||A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|\cos(\Delta\delta_{D})\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (24)
${s_{i}}$ $=\frac{1}{\sqrt{{K_{i}}{K_{-i}}}}\int_{i}|A({{D}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})||A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}})|\sin(\Delta\delta_{D})\text{d}m_{-}^{2}\text{d}m_{+}^{2},$ (25)
where $\Delta\delta_{D}$ is the strong-phase difference between ${D}^{0}$ and ${\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}$ decays to the ${K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ final state. The values used in this measurement for each are taken from Ref. [34]. With these definitions, the parameters in Eq. (4) become $r_{1}\to{r_{K\pi}^{D}}$, $\delta_{1}\to\delta_{K\pi}^{D}$, $R_{1}\to R_{K\pi}=1$, $A_{2}^{2}\to{K_{i}}$, $r_{2}^{2}A_{2}^{2}\to{K_{-i}}$, $R_{2}\cos\delta_{2}\to{c_{i}}$ and $R_{2}\sin\delta_{2}\to{s_{i}}$, such that the fraction of ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays in a region, $i$, is given by,
@@ V.2 Measuring ${{r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}}$ using $D$ decays to $C\!P$-eigenstates
Figures 8 and 9 display the fits for each $C\!P$ tag of the candidates that pass the selection criteria to isolate the $D^{*}\bar{D}\to{\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $D^{*}\bar{D}^{*}\to{\gamma\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanisms, which have the largest yields. The fitted signal yields are corrected for residual backgrounds from ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays to the correct final state, where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair originates from the $e^{+}e^{-}\to D^{*+}D^{*-}$ process, by subtracting the yield in simulation. The corrections are only applied to ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs originating from the $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes and in all cases are less than 7% of the fitted yields. The signal yields after the corrections are displayed in Table 7.
@@
The parameters ${r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}$ and ${r_{K\pi}^{D}}\sin\delta_{K\pi}^{D}$ are determined by simultaneously fitting the reconstructed invariant-mass distribution of both $D$ mesons in 80 categories given by the region of the ${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ phase space and the hypothesized production mechanism. In addition to the five core components discussed previously, two additional background components are included in the fit. The first is from ${{D}^{0}}{{D}^{-}}\to[{{\textit{K}}^{-}}{{\pi}^{+}}]_{{{D}^{0}}},[{K^{0}_{\mathrm{S}}}{{\pi}^{-}}{{\pi}^{0}}]_{{D}^{-}}$ decays, where the ${{D}^{0}}{{D}^{-}}$ pair originates from $e^{+}e^{-}\to{{D}^{*+}}{{D}^{-}}\to{{D}^{0}}{{\pi}^{+}}{{D}^{-}}$ production. These decays are only present in the categories which isolate ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs originating from the $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ process. The selection requirements effectively remove this background in the other categories. The shape and yield of this component are fixed to those determined in simulation. The second background is from ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$ decays, where the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair originates from $e^{+}e^{-}\to D^{*+}D^{*-}$ production. The background is only present in the categories which isolate ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced by the $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ process. In the fit, this background is modeled with the same shape as signal and the yield is fixed to that determined in simulation.
@@ VI Systematic uncertainties
The cross-feed between the production mechanisms is primarily caused by missing energy due to ISR. In events with ISR the effective center-of-mass energy of the $e^{+}e^{-}$ pair is reduced. To model such events the simulation requires knowledge of the cross sections of the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes at the reduced center-of-mass energies, where the relative uncertainties are at most around 20% [36]. The associated systematic uncertainty is determined by repeating the fits which determine the observables using alternate efficiency matrices which are calculated in simulation where the events with ISR are reweighted by ±$\pm 20\%$. Despite this conservative estimate the shifts in the observables are relatively small, for ${r_{K\pi}^{D}}\cos\delta_{K\pi}^{D}$ measured with the $C\!P$ tags (${D}\to{K^{0}_{\mathrm{S}}}{{\pi}^{+}}{{\pi}^{-}}$) it is around 10% (4%) of the statistical uncertainty, whilst for ${r_{K\pi}^{D}}\sin\delta_{K\pi}^{D}$ it is found to be negligible. This systematic uncertainty is fully correlated across all observables.